OpenStudy (lgbasallote):

solve by Laplace transform method \[\Large y^{\prime \prime} + a^2 y =0; \quad y(0) = 0, \; y'(0) = a\]

OpenStudy (lgbasallote):

my solution: \[s^2 y(s) - sy(0) - y^\prime (0) + a^2 y(s) = 0\] \[\implies s^2 y(s) - a + a^2 y(s) = 0\] \[\implies y(s) (s^2 + a^2) - a = 0\] \[\implies y(s) = \frac{a}{s^2 + a^2}\] \[\implies y(t) = \sin at\] am i right?

hartnn (hartnn):

seems absolutely correct...

OpenStudy (lgbasallote):

thanks for confirming